OpenStudy (anonymous):

A man flipped a coin 162 times.the coin landed with heads side up 62 more times than it landed with tails up.how many times did the coin land on heads?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

100

OpenStudy (anonymous):

This is not right.

OpenStudy (across):

\[h+t=162\]\[h=t+62\]\[h-62=t\]\[h+h-62=162\]\[2h-62=162\]\[2h=224\]\[h=112\]

OpenStudy (anonymous):

thanks across

OpenStudy (across):

polpak got it right before I did, though. ;p

OpenStudy (anonymous):

\[h - t = 62\]\[h + t = 162\] Adding the equations we have: \[2h = 224\]\[\implies h = 112\]

OpenStudy (anonymous):

Yeah, but then I accidentally deleted it ;p

OpenStudy (anonymous):

thanks polpak also